OpenStudy (anonymous):

indices= simplify (4^1/2)^3 result equals 8 but can u show how you work it out

OpenStudy (anonymous):

4^(1/2)=(2^2)^(1/2)=2^(2*1/2)=2 hence (4^1/2)^3=2^3=8.@Mars123

OpenStudy (anonymous):

Thank you @Princer_Jones